# thanks: current affiliation: IBM Quantum

### v1

@@ Abstract
Quantum error correction (QEC) codes are necessary to fault-tolerantly operate quantum computers. However, every such code is inherently limited by its inability to detect logical errors. Here, we propose and implement a method that leverages dynamical decoupling (DD) to drastically suppress logical errors. The key to achieving this is to use the logical operators of the QEC code as DD pulses, which we refer to as logical dynamical decoupling (LDD). The resulting hybrid QEC-LDD strategy is in principle capable of handling arbitrary weight errors. We test QEC-LDD using IBM transmon devices and the $[[4,2,2]]$ code, demonstrating performance that significantly exceeds the capabilities of using either this code or DD in isolation. We present a method that allows for the detection of logical errors affecting logically encoded Bell states, which, in this case, arise primarily from crosstalk among physical qubits. Building on this, we experimentally demonstrate high-fidelity entangled logical qubits.
Quantum error correction (QEC) [1, 2, 3, 4] is fundamental to the realization of fault-tolerant quantum computation [5, 6, 7, 8], ensuring the preservation of quantum information undergoing error processes during computation and storage [9, 10]. Numerous successful experimental demonstrations of QEC across various platforms have been reported over the years [11, 12, 13, 14, 15], with the scale and pace accelerating recently towards genuine fault tolerance [16, 17, 18, 19, 18, 20, 21, 22, 23, 24].
Fault-tolerant quantum computing will require the stability of logical qubits over the potentially hours-long timescales of quantum algorithms solving utility-scale problems [25, 26, 27]. Above threshold, over such timescales low-weight physical errors may transform into logical errors of weight higher than any deployed fixed-distance QEC code can handle. This can become a concern even below threshold since long-range spatial and temporal correlations may develop that violate the assumptions underlying fault tolerance theory [28, 29, 30, 31, 32].
@@
Here, we propose and demonstrate a low-cost method that combines a fixed-distance QEC or quantum error detection (QED) code with dynamical decoupling (DD) [39, 40, 41, 42]. This hybrid method, which we call QEC-LDD, can handle arbitrary-weight errors.
Here, LDD stands for ‘logical dynamical decoupling’; the decoupling sequence is implemented using the logical operators of the QEC code as pulses.
We prove a corresponding QEC-LDD theorem which, informally, posits that the most general set of errors acting on an $[[n,k,d]]$ code, namely errors with weights up to $n$, can all either be suppressed using LDD or corrected using the code, for any $d\geq 1$. This seemingly paradoxical result is enabled by virtue of the fact that the LDD sequence suppresses all logical errors as well as some physical errors, thus leaving a much smaller subset of errors for the code to handle, which turn out to all be correctable. Notably, these are not the errors the code was originally designed to correct; indeed, the result holds in particular also for purely error detecting codes. The QEC-LDD theorem is a formal result, in the sense that QEC-LDD, if implemented literally as described in the theorem, is not expected to outperform a pure DD protocol (absent any QEC) on $k$ physical qubits. Nevertheless, the QEC-LDD theorem serves as a useful framework for less ambitious protocols that do not attempt to remove all errors but instead focus on the more physically relevant subset. We discuss and illustrate this conclusion in detail in this work, in the context of a transmon-based demonstration of entangled logical qubits.
Standard, ‘physical’ DD, where DD pulses act not at the logical level but at the physical qubit level, has recently shown great progress, improving the fidelity of storing quantum states [43, 44, 45, 46, 47, 48, 49], circuits [50, 51, 52], and even the performance of entire algorithms [53, 54, 55, 56]. Furthermore, physical DD can be seamlessly combined with fault-tolerant QEC [57, 58] and several recent QEC experiments have used physical DD profitably [18, 21, 19, 59, 20, 60].
@@
However, such pulse sequences can introduce additional errors due to control imperfections in the pulses and due to crosstalk, potentially overshadowing the benefits of DD. To address this challenge, we design our LDD sequences to be robust against such control errors [61, 62] and crosstalk [63, 64, 65, 66, 67], ensuring that the advantages of logical error suppression are not compromised. This robust design allows our LDD implementation to enhance the protection of the code space as intended, effectively targeting both logical and control errors. We demonstrate the practical utility and advantages of this approach using $127$-qubit IBM quantum processors.
@@
The workhorse in our demonstrations is the $[[4,2,2]]$ quantum error detection code, whose two logical qubits we use to prepare logically encoded Bell states. To demonstrate the effectiveness of LDD in suppressing logical errors, we first need a method to detect such errors. However, this code is constrained by its low distance and is incapable of detecting prevalent logical errors, including $ZZ$ crosstalk errors. To overcome this limitation, we design experiments in which, through the use of logically encoded Bell states, we nevertheless unequivocally detect the occurrence of logical errors for a known input logical state. This is then followed by the implementation of various versions of LDD, each employing different subsets of logical operators to construct the sequence. These implementations effectively demonstrate significant suppression of logical errors and substantial improvement in the fidelity of the code space.
@@
The structure of this paper is as follows. In Section I we provide pertinent background on the error model and DD from the perspective of group averaging. We discuss standard Pauli-group DD as well as DD implemented using the stabilizers and or logical operators of a stabilizer code. In Section II we state and prove the QEC-LDD Theorem. We also illustrate the theorem using the example of the $[[4,2,2]]$ code. We then turn to our experimental demonstration of high-fidelity logical Bell states in Section III. This section describes our experimental design and methodology, evidence that both physical DD and error detection with postselection improve logical Bell state fidelity, and finally the evidence that the hybrid QEC-LDD strategy is significantly better.
We conclude in Section IV and provide supporting technical details in the Appendix.
I Background
I.1 Quantum codes and logical errors
An $[[n,k,d]]$ code encodes $k$ logical qubits into $n$ physical qubits with distance $d$ [68]. The weight of an error is the number of physical qubits it affects simultaneously. The highest-weight errors an $[[n,k,d]]$ code can correct or detect are $\lfloor\frac{d-1}{2}\rfloor$ and $d-1$, respectively. A $d=2$ code is a pure quantum error detection (QED) code, and a distance-$d$ QEC code can always be used as a QED code for errors of weight $<d$. The code distance is the minimum number of physical qubits that must experience an error to cause an undetectable logical error, i.e., an error forming a logical operation inside the code space. These logical errors can either be inherently present or result from the accumulation of lower-weight errors over time. For example, in the context of superconducting qubits with fixed-frequency couplers, a prevalent challenge is $ZZ$ crosstalk [69, 70], which inherently introduces weight-two errors that can present as logical errors for distance-$2$ codes.
I.2 Logical dynamical decoupling
@@
Let $\mathcal{P}_{n}=\{\pm P_{j},\pm iP_{j}\}_{j=0}^{4^{n}-1}\subset\mathsf{U}(2^{n})$ be the full Pauli group with $4^{n+1}$ elements generated by $\{I,X,Y,Z\}^{\otimes n}$ and the phase factors {±$\{\pm 1,\pm i\}$. Let $\tilde{\mathcal{P}}_{n}=\mathcal{P}_{n}/\{\pm 1,\pm i\}\subset\mathsf{U}(2^{n})/\mathsf{U}(1)$ be the Abelian factor group with $4^{n}$ elements (here $\mathsf{U}(m)$ is the group of unitary operators on $\mathbb{C}^{m}$).
@@
Consider the most general ‘total decoherence’ system-bath interaction for $n$ qubits:
$H_{SB}=\sum_{j=0}^{4^{n}-1}\alpha_{j}P_{j}\otimes B_{j}.$ (1)
We let $P_{0}=I^{\otimes n}$, the $n$-qubit identity operator. The pure-bath Hamiltonian $P_{0}\otimes H_{B}$ is harmless.
@@
In its simplest form, accounting only for first-order decoupling and ignoring pulse errors, DD theory can be understood as group symmetrization [71, 72]. Choose a decoupling group $\tilde{\mathcal{G}}$ that consists of a set of unitary transformations (pulses) $g_{j}$ acting purely on the system: $\tilde{\mathcal{G}}=\left\{g_{j}\right\}_{j=0}^{\left|\tilde{\mathcal{G}}\right|-1}$, where $g_{0}=P_{0}$. We assume that each such pulse is a unitary operator defined up to a global phase, i.e. $\tilde{\mathcal{G}}\subset\mathsf{U}(2^{n})/\mathsf{U}(1)$; however, to simplify the notation we write, e.g., $g_{0}=P_{0}$ instead of $g_{0}=[P_{0}]$ (the equivalence class of $P_{0}$ up to a global phase). Similarly to $\mathcal{P}_{n}$ and $\tilde{\mathcal{P}}_{n}$, we assume that $\tilde{\mathcal{G}}$ can be written as $\tilde{\mathcal{G}}=\mathcal{G}/\{\pm 1,\pm i\}$ for some group $\mathcal{G}\subset\mathsf{U}(2^{n})$. In particular, we assume that the pulses are instantaneous and cycle over all elements of the group $\tilde{\mathcal{G}}$. This yields the DD pulse sequence:
$U(T)=\prod_{j=0}^{\left|\tilde{\mathcal{G}}\right|-1}g_{j}^{\dagger}f_{\tau}g_{j}=e^{-iT(\langle H_{SB}\rangle_{\mathcal{G}}+P_{0}\otimes H_{B})}+O(T^{2}).$ (2)
@@
Here, $\tau$ is the pulse interval, $T=|\tilde{\mathcal{G}}|\tau$ is the total duration of one repetition of the sequence, $f_{\tau}=e^{-i\tau H_{SB}}$ is the free-evolution unitary, and $\langle H_{SB}\rangle_{\mathcal{G}}$ is the group-symmetrized system-bath Hamiltonian, where
$\langle A\rangle_{\mathcal{G}}\equiv\frac{1}{|\mathcal{G}|}\sum_{g\in\mathcal{G}}g^{\dagger}Ag=\frac{1}{|\tilde{\mathcal{G}}|}\sum_{g\in\tilde{\mathcal{G}}}g^{\dagger}Ag$ (3)
@@
is the projection of the arbitrary bounded system-bath operator $A$ onto the subalgebra of operators that commute with every element of $\mathcal{G}$ (i.e., its commutant). We say that $\tilde{\mathcal{G}}$ decouples $H_{SB}$ (to first order) if $\langle H_{SB}\rangle_{\mathcal{G}}=cI$ where $c$ is a constant, including zero.
@@
For example, $\tilde{\mathcal{G}}=\tilde{\mathcal{P}}_{n}$ decouples an arbitrary $n$-qubit system-bath Hamiltonian since then $\langle H_{SB}\rangle_{\mathcal{G}}=0$, albeit at a sequence time cost of $T=4^{n}\tau$ [40]. To illustrate this, consider $n=1$; the most general system-bath Hamiltonian of a single qubit is $H_{SB}=\sum_{i=0}^{3}\alpha_{i}P_{i}\otimes B_{i}$ with $P_{i}\in\{I,X,Y,Z\}$. If $\tilde{\mathcal{G}}=\{I,X,Y,Z\}$, then $U(T)$ simplifies into the well-known XY4 sequence $U(4\tau)=Yf_{\tau}Xf_{\tau}Yf_{\tau}Xf_{\tau}$ [73], and indeed, $\langle H_{SB}\rangle_{\mathcal{G}}=0$.
@@
Higher-order decoupling sequences are known that achieve suppression up to arbitrary order $q$ [i.e., leaving an $O(T^{q+1})$ unsuppressed error term in $U(T)$[74, 75, 76, 77, 78], but for simplicity, in this work we restrict our attention to first-order sequences. Note that a common misconception is that DD is not effective against purely Markovian noise [79]; however this is not the case, essentially since even in the Markovian limit the bath can have a nonzero correlation time [80, 81, 82, 83].
[…]
Its centralizer is also uniquely defined. However, the group of canonical logical operators is not unique: one can choose arbitrary stabilizers $S_{j}^{X},S_{j}^{Z}\in\mathcal{S}$ ($j=1,\dots,k$) and multiply the generators $\bar{X}_{j},\bar{Z}_{j}$ by those to obtain other choices of the group $\mathcal{L}$.
It is always possible to write the system-bath interaction [Eq. 1] as
$H_{SB}=H_{SB}^{\mathcal{S}}+H_{SB}^{\mathcal{L}}+H_{SB}^{\mathcal{E}},$ (5)

### v2

@@ Abstract
Quantum error correction (QEC) codes are necessary to fault-tolerantly operate quantum computers. However, every such code is inherently limited by its inability to detect logical errors. Here, we propose and implement a method that leverages dynamical decoupling (DD) to drastically suppress logical errors. The key to achieving this is to use the normalizer elements of the QEC code as DD pulses, which we refer to as normalizer dynamical decoupling (NDD). The resulting hybrid QEC-NDD strategy is in principle capable of handling arbitrary weight errors. We test QEC-NDD using IBM transmon devices and the $[[4,2,2]]$ code, demonstrating performance that significantly exceeds the capabilities of using either this code or DD in isolation. We present a method that allows for the detection of logical errors affecting logically encoded Bell states, which, in this case, arise primarily from crosstalk among physical qubits. Building on this, we experimentally demonstrate high-fidelity entangled logical qubits. The fidelities we achieve are beyond-breakeven, i.e., they significantly exceed the corresponding fidelities of unprotected entangled qubits in the same setting.
I Introduction
Quantum error correction (QEC) [1, 2, 3, 4] is fundamental to the realization of fault-tolerant quantum computation [5, 6, 7, 8], ensuring the preservation of quantum information undergoing error processes during computation and storage [9, 10]. Numerous successful experimental demonstrations of QEC across various platforms have been reported over the years [11, 12, 13, 14, 15], with the scale and pace accelerating recently towards genuine fault tolerance [16, 17, 18, 19, 20, 21, 22, 23, 24].
Fault-tolerant quantum computing will require the stability of logical qubits on the long timescales of quantum algorithms that solve utility-scale problems [25, 26, 27]. Above threshold, on such timescales, low-weight physical errors may transform into logical errors, i.e., errors of weight higher than a deployed fixed-distance QEC code can handle. This can become a concern even below threshold, as long-range spatial and temporal correlations may develop that violate the assumptions underlying fault tolerance theory [28, 29, 30, 31, 32].
@@ I Introduction
Here, we propose and demonstrate a low-cost method that combines a fixed-distance QEC or quantum error detection (QED) code with dynamical decoupling (DD) [39, 40, 41, 42]. This hybrid method, which we call QEC-NDD, can handle arbitrary-weight errors. Here, NDD stands for “normalizer dynamical decoupling”; the decoupling sequence is implemented using the normalizer elements of the QEC code as pulses.
@@
Standard, “physical” DD, where DD pulses act not at the logical level but at the physical qubit level, has recently shown great progress, improving the fidelity of storing quantum states [43, 44, 45, 46, 47, 48, 49], circuits [50, 51, 52], and even the performance of entire algorithms [53, 54, 55, 56]. Furthermore, physical DD can be seamlessly combined with fault-tolerant QEC [57, 58] and several recent QEC experiments have used physical DD profitably [18, 21, 19, 59, 20, 60]. However, such pulse sequences can introduce additional errors due to control imperfections in the pulses and due to crosstalk, potentially overshadowing the benefits of DD. To address this challenge, we design our NDD sequences to be robust against such control errors [61, 62] and crosstalk [63, 64, 65, 66, 67], ensuring that the advantages of logical error suppression are not compromised. This robust design allows our NDD implementation to enhance the protection of the code space as intended, effectively targeting both logical and control errors. We demonstrate the practical utility and advantages of this approach using $127$-qubit IBM quantum processors.
@@
The workhorse in our demonstrations is the $[[4,2,2]]$ quantum error detection code, whose two logical qubits we use to prepare logically encoded Bell states. To demonstrate the effectiveness of NDD in suppressing logical errors, we first need a method to detect such errors. However, this code is constrained by its low distance and is incapable of detecting prevalent logical errors, including $ZZ$ crosstalk errors. To overcome this limitation, we design experiments in which, through the use of logically encoded Bell states, we nevertheless unequivocally detect the occurrence of logical errors for a known input logical state. This is then followed by the implementation of various versions of NDD, each employing different subsets of normalizer elements to construct the sequence. These implementations effectively demonstrate significant suppression of logical errors and substantial improvement in the fidelity of the code space.
@@
The structure of this paper is as follows. In Section II we provide pertinent background on the error model and DD from the perspective of group averaging. We discuss standard Pauli group DD, as well as encoded DD, i.e., DD implemented using the stabilizers and/or logical operators of a stabilizer code. In Section III we describe our experimental design and methodology. We then turn to our experimental demonstration of high-fidelity logical Bell states in Section IV. This section describes our experimental design and methodology, evidence that both physical DD and error detection with postselection improve logical Bell state fidelity, and finally, the evidence that the hybrid QEC-NDD strategy significantly outperforms both standalone dynamical decoupling and error detection.
We conclude in Section V and provide supporting technical details in the Appendix.
II Background
In this section, we provide all the theoretical dynamical decoupling background necessary in order to understand the experimental results presented in Section III. In particular, after defining our Hamiltonian error model, we explain how different flavors of encoded DD, i.e., DD implemented using the stabilizers and/or logical operators of a QEC code, suppress the different types of errors that arise in our error model.
II.1 Error model
@@
Let $\mathsf{U}(m)$ denote the group of unitary operators on $\mathbb{C}^{m}$. Let $\mathcal{P}_{n}=\{\pm P_{j},\pm iP_{j}\}_{j=0}^{4^{n}-1}\subset\mathsf{U}(2^{n})$ be the full Pauli group with $4^{n+1}$ elements generated by $\{I,X,Y,Z\}^{\otimes n}$ and the phase factors {±$\{\pm 1,\pm i\}$, where $X\equiv\sigma^{x}$ and likewise for $Y$ and $Z$. There is a natural projection (quotient homomorphism) $\pi:\mathcal{P}_{n}\to\tilde{\mathcal{P}}_{n}$ which takes the Pauli operator and ignores its global phase, yielding a “phase-stripped” Pauli. Correspondingly, let $\tilde{\mathcal{P}}_{n}=\mathcal{P}_{n}/\{\pm 1,\pm i\}\subset\mathsf{U}(2^{n})/\mathsf{U}(1)$ be the Abelian factor group with $4^{n}$ elements, also known as the projective Pauli group.
@@
Consider the most general “total decoherence” system-bath interaction for $n$ qubits:
$H_{SB}=\sum_{{j=1}}^{4^{n}-1}\alpha_{j}P_{j}\otimes B_{j}.$ (1)
In writing this expression we exclude the bath Hamiltonian $I_{S}\otimes B_{0}$ (where $B_{0}=H_{B}$) and select exactly one representative $P_{j}$ from each of the $4^{n}$ equivalence classes $\tilde{\mathcal{P}}_{n}$ such that (s.t.) $P_{0}=I_{S}$ (we use the notation $I_{S}$ to denote the identity operator on the entire system Hilbert space; we use $I$ for a generic identity operator). There is still some freedom in how we order the elements of $\tilde{\mathcal{P}}_{n}$ and which phases we choose. We use this freedom later to organize $\{P_{j}\}_{j=1}^{4^{n}-1}$ in the context of $[[n,k,d]]$ codes.
Our model for $H_{SB}$ is general enough to contain system-only error terms, since any of the bath operators $B_{j}$ may be set equal to the bath identity operator $I_{B}$. Henceforth, when unambiguous, we use $I$ to denote the identity operator regardless of the space on which it acts or its dimension. The joint system-bath free-evolution unitary is $f_{\tau}=e^{-i\tau(H_{SB}+I\otimes H_{B})}$. The adjective “free” refers to the fact that we are leaving out any system-only control Hamiltonian; such control will enter later through DD or QEC. The goal of the latter is, respectively, to suppress or correct the deleterious effects caused by $H_{SB}$.
II.2 Dynamical decoupling and quantum error correction
@@
In its simplest form, accounting only for first-order decoupling and ignoring pulse errors, DD theory can be understood as group symmetrization [68, 69, 70]. Consider a discrete group $\mathcal{G}\subset\mathsf{U}(2^{n})$ with elements $\{g_{j}\}_{j=0}^{|\mathcal{G}|-1}$ representing unitary transformations $g_{j}$ acting purely on the system. We set $g_{0}\equiv g_{|\mathcal{G}|}\equiv I_{S}$ and refer to the corresponding phase-stripped group $\tilde{\mathcal{G}}=\pi(\mathcal{G})=\left\{\tilde{g}_{j}\right\}_{j=0}^{K-1}\subset\mathsf{U}(2^{n})/\mathsf{U}(1)$ as the “decoupling group”, where $K\equiv|\tilde{\mathcal{G}}|=|\mathcal{G}|/4$. This yields the DD cycle
$U(T)$ $=\prod_{j=0}^{K-1}\tilde{g}_{j}^{\dagger}f_{\tau}\tilde{g}_{j}=e^{-iT(\langle H_{SB}\rangle_{\mathcal{G}}+I\otimes H_{B})}+O(T^{2})$ (2)
$=p_{K}f_{\tau}p_{K-1}f_{\tau}\cdots f_{\tau}p_{1}f_{\tau}.$
Here, $\{p_{j}=\tilde{g}_{j}\tilde{g}_{j-1}^{\dagger}\}_{j=1}^{K}$ are the pulses of the corresponding DD pulse sequence, $\tau$ is the pulse interval, and $T=K\tau$ is the total duration of one repetition of the sequence. Note that, equivalently, $\tilde{g}_{j-1}=p_{j-1}\cdots p_{0}$, where $p_{0}\equiv I_{S}$.
@@
$\langle H_{SB}\rangle_{\mathcal{G}}$ is the group-symmetrized system-bath Hamiltonian, where
$\langle A\rangle_{\mathcal{G}}\equiv\frac{1}{|\mathcal{G}|}\sum_{g\in\mathcal{G}}g^{\dagger}Ag=\frac{1}{K}\sum_{\tilde{g}\in\tilde{\mathcal{G}}}\tilde{g}^{\dagger}A\tilde{g}$ (3)
@@
is the projection of the arbitrary bounded system-bath operator $A$ onto the subalgebra of operators that commute with every element of $\mathcal{G}$ (i.e., its commutant). $\langle A\rangle_{\mathcal{G}}$ is also called the group average of $A$ with respect to $\mathcal{G}$.
Definition 1 (Decoupling).
Let $A$ be an arbitrary operator acting on the joint system-bath Hilbert space. $\tilde{\mathcal{G}}$ decouples $A$ (to first order) if $\langle A\rangle_{\mathcal{G}}=cI\otimes B$ where $c$ is a constant, including zero, and $B$ is an arbitrary bath operator.
@@
For example, $\tilde{\mathcal{G}}=\tilde{\mathcal{P}}_{n}$ decouples an arbitrary $n$-qubit system-bath Hamiltonian since then $\langle H_{SB}\rangle_{\mathcal{G}}=0$, albeit at a sequence time cost of $T=4^{n}\tau$ [40]. To illustrate this, consider $n=1$; the most general system-bath Hamiltonian of a single qubit is $H_{SB}=\sum_{{i=1}}^{3}\alpha_{i}P_{i}\otimes B_{i}$ with $P_{i}\in\mathcal{P}_{1}\setminus I$. If $\tilde{\mathcal{G}}={\tilde{\mathcal{P}}_{1}}=\langle X,Z\rangle=\{I,X,Y,Z\}$, then $U(T)$ simplifies into the well-known (universal) XY4 sequence $U(4\tau)=Yf_{\tau}Xf_{\tau}Yf_{\tau}Xf_{\tau}$ [71], and indeed, $\langle H_{SB}\rangle_{\mathcal{G}}=0$. The XX sequence is $U(2\tau)=Xf_{\tau}Xf_{\tau}$, and decouples $Y$ and $Z$, but not $X$ errors. The XY4 sequence can be seen as a concatenation of the XX and ZZ sequences, with the latter defined similarly to XX [72].
@@
Note that a common misconception is that DD is not effective against purely Markovian noise; however, this is not the case, essentially since even in the Markovian limit, the bath can have a nonzero correlation time [73, 74, 75, 76].
In general, decoupling using subgroups of $\tilde{\mathcal{P}}_{n}$ will eliminate parts of $H_{SB}$, presenting an opportunity to selectively combine DD with QEC [77, 57]. See Appendix A for pertinent background on QEC codes.
Let $\mathcal{S}$ be a stabilizer group specifying an $[[n,k,d]]$ stabilizer code. The group of logical operators is $\mathcal{N}(\mathcal{S})/\mathcal{S}$, where $\mathcal{N}(\mathcal{S})$ is the normalizer of $\mathcal{S}$ in $\mathcal{P}_{n}$ (which here coincides with the centralizer, i.e., the group of Pauli operators commuting with all elements of $\mathcal{S}$). We can pick a canonical set of generators of the stabilizer group and a canonical group of logical operators $\mathcal{L}$ [^1] That is, any transversal for $\mathcal{S}$ in $\mathcal{N}(\mathcal{S})$ satisfying $\pi^{-1}(\pi(\mathcal{L}))=\mathcal{L}$ can be picked as $\mathcal{L}$, i.e., any subgroup of $\mathcal{N}(\mathcal{S})$ with exactly one representative from each coset $L\mathcal{S}\subset\mathcal{N}(\mathcal{S})$ s.t. $iI\in\mathcal{L}$. Any such $\mathcal{L}$ is isomorphic to $\mathcal{P}_{k}$ and can be thought of as a group of Paulis acting on logical qubits.
We can use the remaining freedom mentioned above to reorder $\{P_{j}\}_{j=0}^{4^{n}-1}$ and pick the phases of those operators to ensure that
  • $P_{0}=I_{S}$
  • $\mathcal{S}_{\star}=\{P_{1},\dots,P_{2^{n-k}-1}\}=\mathcal{S}\setminus\{I_{S}\}$ (these are non-identity stabilizers acting trivially on the code space)
  • =$\mathcal{L}_{\star}=\{P_{2^{n-k}},\dots,P_{2^{n+k}-1}\}\subset\mathcal{N}(\mathcal{S})\setminus\pi^{-1}(\tilde{\mathcal{S}})$ (these manifest as undetectable logical errors when acting on the code space)
  • $\mathcal{D}_{\star}=\{P_{2^{n+k}},\dots,P_{4^{n}-1}\}\subset\mathcal{P}_{n}\setminus\mathcal{N}(\mathcal{S})$ (these are detectable errors).
We chose the phases of the representatives of $\tilde{\mathcal{S}}$ to ensure they are actually in $\mathcal{S}$ (and not different from elements of $\mathcal{S}$ by a phase). This represents $\{P_{j}\}_{j=0}^{4^{n}-1}$ as a disjoint union
$\{P_{j}\}_{j=0}^{4^{n}-1}=\{I_{S}\}\dot{\cup}\mathcal{S}_{\star}\dot{\cup}\mathcal{L}_{\star}\dot{\cup}\mathcal{D}_{\star}.$ (4)
The system Hilbert space $\mathcal{H}$ can be split as
$\mathcal{H}=\mathcal{H}_{\text{stab}}\otimes\mathcal{H}_{\text{log}},$ (5)
where $\mathcal{S}$ acts on the first component and $\mathcal{L}$ acts on the second, and where $\dim(\mathcal{H}_{\text{stab}})=2^{n-k}$ and $\dim(\mathcal{H}_{\text{log}})=2^{k}$, corresponding to the $n-k$ (virtual) syndrome qubits and $k$ logical qubits, respectively.
With respect to this virtual tensor product decomposition induced by the syndrome measurement [78], every Pauli operator $E\in\mathcal{S}_{\star}$, $\mathcal{L}_{\star}$, or $\mathcal{D}_{\star}$ can be written as:
  • $E=P\otimes I\in\mathcal{S}_{\star}$ with $P\in\mathcal{P}_{n-k}$ (a stabilizer with trivial action on the code space)
  • $E=P\otimes Q\in\mathcal{L}_{\star}$ with $[P,\mathcal{S}_{\star}]=0$ (a logical operator manifesting as an error if it appears in $H_{SB}$).
  • $E=P\otimes Q\in\mathcal{D}_{\star}$ with $\{P,S_{i}\}=0$ for one or more $S_{i}\in\mathcal{S}_{\star}$ (a detectable error).
Correspondingly, it is always possible to write the system-bath interaction [Eq. 1] as
$H_{SB}$ $=H_{SB}^{\mathcal{S}_{\star}}+H_{SB}^{\mathcal{L}_{\star}}+H_{SB}^{\mathcal{D}_{\star}}$ (6a)
$H_{SB}^{\mathcal{S}_{\star}}\!\!=\sum_{i=1}^{2^{n-k}-1}S_{i}\otimes B^{\text{stab}}_{i}\ ,\quad S_{i}\in\mathcal{S}_{\star}$ (6b)
$H_{SB}^{\mathcal{L}_{\star}}=\sum_{i=1}^{2^{n-k}(4^{k}-1)}L_{i}\otimes B^{\text{log}}_{i}\ ,\quad L_{i}\in\mathcal{L}_{\star}$ (6c)
$H_{SB}^{\mathcal{D}_{\star}}=\sum_{i=1}^{4^{n}-2^{n+k}}D_{i}\otimes B^{\text{det}}_{i}\ ,\quad D_{i}\in\mathcal{D}_{\star},$ (6d)
where $\{B^{\text{stab}}_{i},B^{\text{log}}_{i},B^{\text{det}}_{i}\}$ are bath operators with units of energy.